One of our LNS routers in London Telehouse is experiencing performance issues. The affected device is being rebooted to clear the issue. Customers with DSL sessions terminated on the device will have loss of service until the problem is rectified.
Update 07:35 - Device was rebooted, customer DSL sessions are now back online. Engineers are continuing to monitor.
28th June 2017
London Telehouse Optical Wavelength Maintenance, scheduled 1 day ago
We will be introducing new optical wavelengths into our London Telehouse sites on Weds 28th June. This work should not be service affecting but due to engineers working within core network rack cabinets there is minor risk of disruption.